SpaceRef

SpaceRef


NOAA SATOPS Morning Report: Tuesday, December 13, 2011

Status Report From: NOAA Office of Satellite Operations
Posted: Tuesday, December 13, 2011

Customers or users who want to call to ask about problems or report outages can use the following contacts:

- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 347

GOES
All GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
OPS54 DOY347 02:01z Cat 2 data loss of 6173 s/c seconds due to a POGO ground station problem.

All other DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety

ESPC Reported data outage and delays

A communications line between NESDIS and the NWS Telecommunications Gateway failed at around on December 12 at approximately 1940 UTC. This is affecting some NESDIS products that are distributed via the satepsdist1e product server. The problem has been isolated to Verizon and they are working on a resolution. Impacted products include: DMSP SSMI/S F-15, 16,17 and 18 data, Cloud North and South data, HRPT data from Monterey, MCST and Windsat data.

ESPC is not receiving MET9 data due to a DS-3 comm link from Wallops to Darmstadt Under investigated.

The following GOES-west GINI images were lost due to a NSOF equipment problem: 1000z, 1015z, 1030z, 1045z, and 1100z.

Questions or comments to Ron.Mahmot@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks